which is an example of distibutive property ?
A. (5-2)+3=(-2+5)+3
B. 5(2-3)=5X2-5X3

Answer by Theo(13342)

distributive property is:
a(b+c) = ab + ac
that would be selection b.
5(2-3) = 5(2) - 5(3)
selection a would be an example of the commutative property i believe.
that one says that a+b = b+a which is what you see within the parentheses.
selection a says:
(5-2) + 3 = (-2+5) + 3
the position of the constants within the parentheses have been reversed which is the commutative property in action.